OpenStudy (anonymous):

evaluate the series 1+2+4+8+...+a10

OpenStudy (mathmath333):

what is \(a10\)?

OpenStudy (mathmath333):

i assume it upto 10 terms

OpenStudy (mathmath333):

its Geometric series with common ratio(r)=2 r=2nd term/1st term=3rd term/2nd term......=2/1=4/2.......=2

OpenStudy (mathmath333):

Sum of a_10 terms= a(r^n-1)/((r-1) where n = number of terms and a= first term

OpenStudy (mathmath333):

a=1 r=2 n=10